Express the following expression in simplest radical form. All variables represent positive real numbers.
SQRT of (8y^3)/49x

Answer by fractalier(6550)

From sqrt(8y^3 / 49x), we will need to rationalize the denominator by multiplying by sqrt(x) / sqrt(x)...this will give us a perfect square in the denominator...
sqrt(8y^3 / 49x) =
sqrt(8y^3 / 49x) * [sqrt(x) / sqrt(x)] =
sqrt(8xy^3 / 49x^2)
We can pull perfect squares out of the radical from the numerator and denominator, giving us the answer...
sqrt(8xy^3 / 49x^2) =
(2y / 7x) * sqrt(2xy)